# On weight and variance uncertainty in neural networks for regression tasks

### v1

@@ Abstract
We consider the problem of weight uncertainty proposed by [Blundell et al. (2015). Weight uncertainty in neural network. In International conference on machine learning, 1613-1622, PMLR.] in neural networks (NNs) specialized for regression tasks. We further investigate the effect of variance uncertainty in their model. We show that including the variance uncertainty can improve the prediction performance of the Bayesian NN. Variance uncertainty enhances the generalization of the model by considering the posterior distribution over the variance parameter. We examine the generalization ability of the proposed model using a function approximation example and further illustrate it with the riboflavin genetic data set. We explore fully connected dense networks and dropout NNs with Gaussian and spike-and-slab priors, respectively, for the network weights.
Keywords: Bayesian neural networks, Variational Bayes, Kullback-Leibler divergence, posterior distribution, Regression task, variance uncertainty.
@@ 1 Introduction
Bayesian Neural Networks (BNNs) have been introduced and comprehensively discussed by many authors (among others see [14, 2]). BNNs are suitable for modeling uncertainty by considering values of the parameters that might not be learned by the available data. This is achieved by randomizing unknown parameters, such as weights and biases while incorporating prior knowledge. Such an approach naturally regularizes the model and helps prevent overfitting, a common challenge in neural networks (NNs). The BNNs can also be considered as a suitable and more reliable alternative to ensemble learning methods [15], including bagging and boosting of the neural networks.
@@
Another approach for modeling uncertainty in weights of a NN is considered in [7]. They modeled the variational approximation of the weights as the multiplication of a fixed hyperparameter (which might be optimized) by a bernoulli random varaible for modeling the droupout mechanism. There are two main differences between the method considered in [7] and that in [4]. The first difference is that no statistical distribution is considered for modelling uncertainty in weights around the hyperparameter in the method proposed in [7]. The second difference is the approach for modeling the dropout mechanishm, since the dropout mechinism was modeled by considering the slab-and-spike prior for the weights of the NN in [4].
@@ 2 Bayesian Neural Networks
In this section, our goal is to describe BNNs and their challenges in practical applications. Before examining BNNs, we briefly describe NNs from a statistical perspective. A NN aims to approximate the unknown function $y=\phi(x)$ where $x$ and $y$ are the input and the output vectors, respectively. From a statistical point of view, given the sample points $x,y$, the objective function of NN is the log-likelihood function of weights and biases, $W$, and extra parameters $\rho$, of the model, $\log L(W,\rho|x,y)$ which is aimed to be maximized to learn the parameters of the network
$\widehat{(W,\rho)}_{\text{\rm MLE}}=\arg\max_{(W,\rho)}\log L(W,\rho\mid x,y).$
The likelihood function $L(W,\rho\mid x,y)$ depends on the specific problem being addressed:
  • In regression tasks, the likelihood function is typically modeled by a Gaussian density of $y$ with a mean provided by the network given the input data set $x$ and an unknown variance $\sigma^{2}$, and the negative log-likelihood corresponds to the sum of squared errors (SSE).
@@
In a BNN, a prior distribution $p(W,\rho)$ is considered for the parameters of the model [14], and our goal is to find a posterior distribution of all parameters $p(W,\rho\mid x,y)$. Using the Bayes’ formula, the posterior density is obtained as
$p(W,\rho\mid x,y)=\frac{L(W,\rho\mid x,y)\,p(W,\rho)}{p(x,y)},$
where $p(x,y)$ is the marginal distribution of the data, given by
$p(x,y)=\int L(W,\rho\mid x,y)\,p(W,\rho)\,d(W,\rho).$
@@
$\widehat{(W,\rho)}_{\text{\rm MAP}}=\arg\max_{(W,\rho)}\log L(W,\rho\mid x,y)+\log p(W,\rho).$
@@
$p(y^{*}\mid x,y,x^{*})=\int_{\tilde{\Theta}}L(W,\rho\mid x^{*},y^{*})\,p(W,\rho\mid x,y)\,d(W,\rho).$
In practice, we face intractable integrals to solve $p(x,y)$; therefore, we resort to approximation methods. In the following, we introduce VB as a well-known approximation technique.
@@ 3 Variational Bayes
The VB is a method for finding an approximate distribution $q(\theta)$ of the posterior distribution $p(\theta|x)$, by minimizing the Kullback-Leibler divergence ${\rm KL}\left[q(\theta)||p(\theta|x)\right]$ as a measure of closeness [18]. In this section, we briefly review the basic elements of the VB method.
[…]
In the VB method, we assume the parameter vector $\theta$ is divided into $M$ partitions $\{\theta_{1},\ldots,\theta_{M}\}$ and we want to approximate $p(\theta|x)$ by
$q(\theta)=\prod_{j=1}^{M}q_{j}(\theta_{j}).$
where $q_{j}(.)$ is the variational posterior over the vector of posterior parameters $\theta_{j}$. Hence, each $\theta_{j}$ can be a vector of parameters. The best VB approximation $q^{\ast}(\cdot)$ is then obtained as
$q^{\ast}=\arg\min_{q}{\rm KL}\left[q(\theta)||p(\theta|x)\right],$
@@
${\rm KL}\left[q(\theta)||p(\theta|x)\right]=\int q(\theta)\log\frac{q(\theta)}{p(\theta|x)}d\theta=\log p(x)-\int q(\theta)\log\frac{p(x,\theta)}{q(\theta)}d\theta.$ (1)
Let $\eta=\{\eta_{1},\ldots,\eta_{M}\}$ be the partitions of hyper-parameter such that each $\eta_{j}$ be the variational parameter of $\theta_{j}$. In the fixed-form-VB (FFVB), we consider an assumed density function for each $q(\theta_{j}|\eta_{j})$, with unknown $\eta_{j}$, and we aim to determine the hyper-parameters $\eta$ such that ${\rm KL}\left[q(\theta|\eta)||p(\theta|x)\right]$ is minimized.
It is clear from (1) that minimizing KL is equivalent to maximizing the evidence lower bound (ELBO) with respect to $\eta$, defined as
${\rm ELBO}=\int q(\theta|\eta)\log\frac{p(x,\theta)}{q(\theta|\eta)}d\theta$
The ELBO is equal to $\log p(x)$ when the KL divergence is zero, meaning a perfect fit. The closer the value of ${\rm ELBO}[q(\theta)]$ to $\log p(x)$ (greater values), the fit is better. Maximizing the ELBO is equivalent to minimizing its negative. It follows that
$\eta^{*}=\inf_{\eta}\int q(\theta|\eta)\log\frac{q(\theta|\eta)}{p(x|\theta)p(\theta)}\,d\theta=\inf_{\eta}\mathbb{E}_{q(\theta|\eta)}\left[\log q(\theta|\eta)-\log p(x|\theta)-\log p(\theta)\right],$
where $\eta^{*}$ represents the optimal parameter estimate for maximizing the ELBO.
We define a loss function to be minimized. Assume that $\epsilon$ is a random variable with a probability density function $q(\epsilon)$, and let $f$ be a function. In this case, as shown in [4], if
$q(\epsilon)d\epsilon=q(\theta|\eta)d\theta,$ (2)
@@
$\frac{\partial}{\partial\eta}\mathbb{E}_{q(\theta|\eta)}[f(\theta,\eta)]=\mathbb{E}_{q(\epsilon)}\left[\frac{\partial f(\theta,\eta)}{\partial\eta}\right].$ (3)
Equation (3) shows that if we can find a transformation for $\theta$ such that it depends on $\epsilon$ and satisfies the condition (2), then samples of $\epsilon$ can be used to generate samples of $\theta$. This approach focuses on finding $\eta^{*}$, as defined above. Therefore, we define
$f(\theta,\eta)=\log q(\theta|\eta)-\log p(x|\theta)-\log p(\theta).$ (4)
Note that the expectation of $f(\theta,\eta)$ is the negative of the ELBO and the gradient of $f(\theta,\eta)$ provides an unbiased estimate of the gradient defined in equation (3). In addition, increasing the number of samples of $\epsilon$ reduces the variance of the estimated gradient. This method utilizes equation (4) as an objective function to determine $\eta^{*}$.
@@ 4 Variational Bayes Regression Network with Fixed Variance
Following the approach described in [4], the variance of the likelihood function in a regression problem can be fixed to a constant value. In regression problems, assuming independency among the output features (i.e., the values of the vector $y_{i}$), we have
$y_{i}=\phi(x_{i};W)+\epsilon_{i},\quad\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2}_{0}I_{q}),$
where $I_{q}$ is the identity matrix, $x_{i}\in\mathbb{R}^{p}$ represents the $i$-th input variable, $y_{i}\in\mathbb{R}^{q}$ denotes the $i$-th output variable, $i=1,...,n$, and $\sigma^{2}_{0}$ is a fixed known variance. Moreover, $\phi(x_{i};W)$ represents the output of a NN with an arbitrary number of layers. In this case, we have the conditional distribution
$y_{i}|x_{i},W\sim\mathcal{N}(\phi(x_{i};W),\sigma^{2}_{0}I_{q}).$
If $(x_{i},y_{i})$ is the $i$th observation, then the likelihood function, assuming independent observations, is given by
$L(W\mid y,x)$ $=p(y_{1},\dots,y_{n}\mid x_{1},\dots,x_{n},W)$
$=\prod_{i=1}^{n}\frac{1}{\sqrt{2\pi\sigma_{0}^{2}}}\exp\left\{-\frac{1}{2\sigma_{0}^{2}}(y_{i}-\phi(x_{i};W))^{T}(y_{i}-\phi(x_{i};W))\right\}$
@@
$f(W,\eta)=\log q(W|\eta)-\log L(W\mid y,x)-\log p(W).$ (5)
with
$q(W|\eta)=\mathcal{N}(W;\mu_{w},\log(1+\exp(\rho_{w}))),$
and $\eta=(\mu_{w},\rho_{w})$, seeking the optimum value of $\eta$. We denote this method by VBNET-FIXED throughout the remaining of the paper. However, this approach fails to effectively capture uncertainty in the variance parameter. In the forthcoming section, we will discuss how to introduce uncertainty in the variance by using parametrization, so that the model can be generalized by considering the posterior distribution over the variance parameter.

### v2

@@ Abstract
We investigate the problem of weight uncertainty originally proposed by [Blundell et al. (2015). Weight uncertainty in neural networks. In International conference on machine learning, 1613-1622, PMLR.] in the context of neural networks designed for regression tasks, and we extend their framework by incorporating variance uncertainty into the model. Our analysis demonstrates that explicitly modeling uncertainty in the variance parameter can significantly enhance the predictive performance of Bayesian neural networks. By considering a full posterior distribution over the variance, the model achieves improved generalization compared to approaches that treat variance as fixed or deterministic. We evaluate the generalization capability of our proposed approach through a function approximation example and further validate it on the riboflavin genetic dataset. Our exploration encompasses both fully connected dense networks and dropout neural networks, employing Gaussian and spike-and-slab priors respectively for the network weights, providing a comprehensive assessment of how variance uncertainty affects model performance across different architectural choices.
Keywords: Bayesian neural networks, Kullback-Leibler divergence, Posterior distribution, Regression task, Variance uncertainty, Variational Bayes.
@@ 1 Introduction
Bayesian Neural Networks (BNNs) have been introduced and comprehensively discussed by many authors (among others see [16, 3]). BNNs are suitable for modeling uncertainty by considering values of the parameters that might not be learned by the available data. This is achieved by randomizing unknown parameters, such as weights and biases while incorporating prior knowledge. Such an approach naturally regularizes the model and helps prevent overfitting, a common challenge in neural networks. Recently, other approaches such as heuristic shrinkage penalized methods have also been proposed to address the regression problems in deep neural networks [2]. The BNNs can also be considered as a suitable and more reliable alternative to ensemble learning methods [17], including bagging and boosting of the neural networks.
@@
Another approach for modeling uncertainty in weights of a NN is considered in [8]. They modeled the variational approximation of the weights as the multiplication of a fixed hyperparameter (which might be optimized) by a bernoulli random varaible for modeling the dropout mechanism. There are two main differences between the method considered in [8] and that in [5]. The first difference is that no statistical distribution is considered for modelling uncertainty in weights around the hyperparameter in the method proposed in [8]. The second difference is the approach for modeling the dropout mechanism, since the dropout mechanism was modeled by considering the slab-and-spike prior for the weights of the NN in [5].
@@
While weight uncertainty has been extensively studied, treating the likelihood variance as a random variable in BNNs offers distinct advantages that are often overlooked. We distinguish our approach from heteroscedastic regression models (e.g., [15]), where the network predicts a data-dependent variance $\sigma(x)$ for each input. In contrast, our work focuses on the epistemic uncertainty of the global observation variance. This is particularly important in limited-data problems, where the true variance is unknown and assuming a fixed or point-estimated variance can lead to overconfident predictions. Furthermore, unlike classical Bayesian models which typically employ an Inverse-Gamma conjugate prior for the variance, such closed-form updates are intractable in deep neural networks. Our proposed method extends the Bayes-by-Backprop framework to the likelihood variance, enabling joint optimization of weights and variance parameters via stochastic gradient descent without requiring conjugacy. Finally, compared to optimizing $\sigma^{2}$ as a point estimate, marginalizing over the variance posterior introduces heavy-tailed behavior in the predictive distribution. This results in a model that is more robust to outliers and provides better prediction intervals, as demonstrated by our coverage probability results in Section 6.
@@ 2 Bayesian Neural Networks
In this section, our goal is to describe BNNs and their challenges in practical applications. Before examining BNNs, we briefly describe NNs from a statistical perspective. A NN aims to approximate the unknown function $\mathbf{y}={\phi(\mathbf{x})}$ where $\phi(\mathbf{x})$} and $\mathbf{y}$ are the input and the output vectors, respectively. From a statistical point of view, given the sample points $\mathbf{x},\mathbf{y}$, the objective function of NN is the log-likelihood function of weights and biases, $\mathbf{W}$, and extra parameters $\rho$, of the model, $\log L(\mathbf{W},\rho|\mathbf{x},\mathbf{y})$ which is aimed to be maximized to learn the parameters of the network
$\widehat{(\mathbf{W},\rho)}_{\text{\rm MLE}}=\arg\max_{(\mathbf{W},\rho)}\log L(\mathbf{W},\rho\mid\mathbf{x},\mathbf{y}).$
The likelihood function $L(\mathbf{W},\rho\mid\mathbf{x},\mathbf{y})$ depends on the specific problem being addressed:
  • In regression tasks, the likelihood function is typically modeled by a Gaussian density of $\mathbf{y}$ with a mean provided by the network given the input data set $\mathbf{x}$ and an unknown variance $\sigma^{2}$, and the negative log-likelihood corresponds to the sum of squared errors (SSE).
@@
In a BNN, a prior distribution $p(\mathbf{W},\rho)$ is considered for the parameters of the model [16], and our goal is to find a posterior distribution of all parameters $p(\mathbf{W},\rho\mid\mathbf{x},\mathbf{y})$. Using the Bayes’ formula, the posterior density is obtained as
$p(\mathbf{W},\rho\mid\mathbf{x},\mathbf{y})=\frac{L(\mathbf{W},\rho\mid\mathbf{x},\mathbf{y})\,p(\mathbf{W},\rho)}{p(\mathbf{x},\mathbf{y})},$
where $p(\mathbf{x},\mathbf{y})$ is the marginal distribution of the data, given by
$p(x,y)=\int L(\mathbf{W},\rho\mid\mathbf{x},\mathbf{y})\,p(\mathbf{W},\rho)\,d(\mathbf{W},\rho).$
@@
$\widehat{(\mathbf{W},\rho)}_{\text{\rm MAP}}=\arg\max_{(\mathbf{W},\rho)}\log L(\mathbf{W},\rho\mid\mathbf{x},\mathbf{y})+\log p(\mathbf{W},\rho).$
@@
$p(\mathbf{y^{*}}\mid\mathbf{x},\mathbf{y},\mathbf{x^{*}})=\int_{\tilde{\boldsymbol{\theta}}}L(\mathbf{W},\rho\mid\mathbf{x^{*}},\mathbf{y^{*}})\,p(\mathbf{W},\rho\mid\mathbf{x},\mathbf{y})\,d(\mathbf{W},\rho).$
In practice, we face intractable integrals to solve $p(\mathbf{x},\mathbf{y})$; therefore, we resort to approximation methods. In the following, we introduce VB as a well-known approximation technique.
@@ 3 Variational Bayes
The VB is a method for finding an approximate distribution $q(\boldsymbol{\theta})$ of the posterior distribution $p(\boldsymbol{\theta}|\mathbf{x})$, by minimizing the Kullback-Leibler divergence ${\rm KL}\left[q(\boldsymbol{\theta})||p(\boldsymbol{\theta}|\mathbf{x})\right]$ as a measure of closeness [tr23]. In this section, we briefly review the basic elements of the VB method.
Let $\mathbf{x}$ be a vector of observed data, and $\boldsymbol{\theta}$ be a parameter vector with joint distribution $p(\mathbf{x},\boldsymbol{\theta})$. In the Bayesian inference framework, the inference about $\boldsymbol{\theta}$ is done based on the posterior distribution $p(\boldsymbol{\theta}|\mathbf{x})=p(\mathbf{x},\boldsymbol{\theta})/p(\mathbf{x})$ where $p(\mathbf{x})=\int p(\mathbf{x},\boldsymbol{\theta})d\boldsymbol{\theta}$.
In the VB method, we assume the parameter vector $\boldsymbol{\theta}$ is divided into $M$ partitions $\{\boldsymbol{\theta}_{1},\ldots,\boldsymbol{\theta}_{M}\}$ and we want to approximate $p(\boldsymbol{\theta}|\mathbf{x})$ by
$q(\boldsymbol{\theta})=\prod_{j=1}^{M}q_{j}(\boldsymbol{\theta}_{j}),$
where $q_{j}(.)$ is the variational posterior over the vector of posterior parameters $\boldsymbol{\theta}_{j}$. Hence, each $\boldsymbol{\theta}_{j}$ can be a vector of parameters. The best VB approximation $q^{\ast}(\cdot)$ is then obtained as
$q^{\ast}=\arg\min_{q}{\rm KL}\left[q(\boldsymbol{\theta})||p(\boldsymbol{\theta}|\mathbf{x})\right],$
@@
${\rm KL}\left[q(\boldsymbol{\theta})||p(\boldsymbol{\theta}|\mathbf{x})\right]=\int q(\boldsymbol{\theta})\log\frac{q(\boldsymbol{\theta})}{p(\boldsymbol{\theta}|\mathbf{x})}d\boldsymbol{\theta}=\log p(\mathbf{x})-\int q(\boldsymbol{\theta})\log\frac{p(\mathbf{x},\boldsymbol{\theta})}{q(\boldsymbol{\theta})}d\boldsymbol{\theta}.$ (1)
Let $\boldsymbol{\eta}=\{\boldsymbol{\eta}_{1},\ldots,\boldsymbol{\eta}_{M}\}$ be the partitions of hyper-parameter such that each $\boldsymbol{\eta}_{j}$ be the variational parameter of $\boldsymbol{\theta}_{j}$. In the fixed-form-VB (FFVB), we consider an assumed density function for each $q(\boldsymbol{\theta}_{j}|\boldsymbol{\eta}_{j})$, with unknown $\boldsymbol{\eta}_{j}$, and we aim to determine the hyper-parameters $\boldsymbol{\eta}$ such that ${\rm KL}\left[q(\boldsymbol{\theta}|\boldsymbol{\eta})||p(\boldsymbol{\theta}|x)\right]$ is minimized.
It is clear from (1) that minimizing KL is equivalent to maximizing the evidence lower bound (ELBO) with respect to $\boldsymbol{\eta}$, defined as
${\rm ELBO}=\int q(\boldsymbol{\theta}|\boldsymbol{\eta})\log\frac{p(\mathbf{x},\boldsymbol{\theta})}{q(\boldsymbol{\theta}|\boldsymbol{\eta})}d\boldsymbol{\theta}.$
The ELBO is equal to $\log p(x)$ when the KL divergence is zero, meaning a perfect fit. The closer the value of ${\rm ELBO}[q(\boldsymbol{\theta})]$ to $\log p(\mathbf{x})$ (greater values), the fit is better. Maximizing the ELBO is equivalent to minimizing its negative. It follows that
$\boldsymbol{\eta}^{*}=\inf_{\boldsymbol{\eta}}\int q(\boldsymbol{\theta}|\boldsymbol{\eta})\log\frac{q(\boldsymbol{\theta}|\boldsymbol{\eta})}{p(\mathbf{x}|\boldsymbol{\theta})p(\boldsymbol{\theta})}\,d\boldsymbol{\theta}=\inf_{\boldsymbol{\eta}}\mathbb{E}_{q(\boldsymbol{\theta}|\boldsymbol{\eta})}\left[\log q(\boldsymbol{\theta}|\boldsymbol{\eta})-\log p(\mathbf{x}|\boldsymbol{\theta})-\log p(\boldsymbol{\theta})\right],$
where $\boldsymbol{\eta}^{*}$ represents the optimal parameter estimate for maximizing the ELBO.
We define a loss function to be minimized. Assume that $\varepsilon$ is a random variable with a probability density function $q(\varepsilon)$, and let $f$ be a function. In this case, as shown in [5], if
$q(\varepsilon)d\varepsilon=q(\boldsymbol{\theta}|\boldsymbol{\eta})d\boldsymbol{\theta},$ (2)
@@
$\frac{\partial}{\partial\boldsymbol{\eta}}\mathbb{E}_{q(\boldsymbol{\theta}|\boldsymbol{\eta})}[f(\boldsymbol{\theta},\boldsymbol{\eta})]=\mathbb{E}_{q(\varepsilon)}\left[\frac{\partial f(\boldsymbol{\theta},\boldsymbol{\eta})}{\partial\boldsymbol{\eta}}\right].$ (3)
Equation (3) shows that if we can find a transformation for $\boldsymbol{\theta}$ such that it depends on $\varepsilon$ and satisfies the condition (2), then samples of $\varepsilon$ can be used to generate samples of $\boldsymbol{\theta}$. This approach focuses on finding $\boldsymbol{\eta}^{*}$, as defined above. Therefore, we define
$f(\boldsymbol{\theta},\boldsymbol{\eta})=\log q(\boldsymbol{\theta}|\boldsymbol{\eta})-\log p(\mathbf{x}|\boldsymbol{\theta})-\log p(\boldsymbol{\theta}).$ (4)
Note that the expectation of $f(\boldsymbol{\theta},\boldsymbol{\eta})$ is the negative of the ELBO and the gradient of $f(\boldsymbol{\theta},\boldsymbol{\eta})$ provides an unbiased estimate of the gradient defined in equation (3). In addition, increasing the number of samples of $\varepsilon$ reduces the variance of the estimated gradient. This method utilizes equation (4) as an objective function to determine $\boldsymbol{\eta}^{*}$.
@@ 4 Variational Bayes Regression Network with Fixed Variance
Following the approach described in [5], the variance of the likelihood function in a regression problem can be fixed to a constant value. In regression problems, assuming independency among the output features (i.e., the values of the vector $\mathbf{y}_{i}$), we have
$\mathbf{y}_{i}=\phi(\mathbf{x}_{i};\mathbf{W})+\varepsilon_{i},\quad\varepsilon_{i}\sim\mathcal{N}(0,\sigma^{2}_{0}\mathbf{I_{q}}),$
where $\mathbf{I_{q}}$ is the identity matrix, $\mathbf{x}_{i}\in\mathbb{R}^{p}$ represents the $i$-th input variable, $\mathbf{y}_{i}\in\mathbb{R}^{q}$ denotes the $i$-th output variable, i=1,,ni=1,...,n, and $\sigma^{2}_{0}$ is a fixed known variance. Moreover, $\phi(\mathbf{x}_{i};\mathbf{W)}$ represents the output of a NN with an arbitrary number of layers. In this case, we have the conditional distribution
$\mathbf{y}_{i}|\mathbf{x}_{i},\mathbf{W}\sim\mathcal{N}(\phi(\mathbf{x}_{i};\mathbf{W}),\sigma^{2}_{0}\mathbf{I_{q}}).$
If $(\mathbf{x}_{i},\mathbf{y}_{i})$ is the $i$th observation, then the likelihood function, assuming independent observations, is given by
$L(\mathbf{W}\mid\mathbf{y},\mathbf{x})$ $=p(\mathbf{y}_{1},\dots,\mathbf{y}_{n}\mid\mathbf{x}_{1},\dots,\mathbf{x}_{n},\mathbf{W})$
$=\prod_{i=1}^{n}\frac{1}{\sqrt{{\color[rgb]{0,0,0}(2\pi\sigma_{0}^{2})^{q}}}}\exp\left\{-\frac{1}{2\sigma_{0}^{2}}(\mathbf{y}_{i}-\phi(\mathbf{x}_{i};\mathbf{W}))^{T}(\mathbf{y}_{i}-\phi(\mathbf{x}_{i};\mathbf{W}))\right\}.$
@@
$f(\mathbf{W},\boldsymbol{\eta})=\log q(\mathbf{W}|\boldsymbol{\eta})-\log L(\mathbf{W}\mid\mathbf{y},\mathbf{x})-\log p(\mathbf{W}),$ (5)
where $q(\mathbf{W}|\boldsymbol{\eta})=\mathcal{N}(\mathbf{W};\boldsymbol{\mu}_{w},\operatorname{diag}(\boldsymbol{\sigma}_{w}^{2}))$ with $\boldsymbol{\sigma}_{w}=\log(1+\exp(\boldsymbol{\rho}_{w}))$, and $\boldsymbol{\eta}=(\boldsymbol{\mu}_{w},{\boldsymbol{\rho}_{w}})$, seeking the optimum value of $\boldsymbol{\eta}$. We denote this method by VBNET-FIXED throughout the remaining of the paper. However, this approach fails to effectively capture uncertainty in the variance parameter. In the forthcoming section, we will discuss how to introduce uncertainty in the variance by using parametrization, so that the model can be generalized by considering the posterior distribution over the variance parameter.